OpenStudy (anonymous):

A sequence is defined by the formula f(n + 1) = f(n) – 3. If f(4) = 22, what is f(1)?

OpenStudy (anonymous):

I figured out the answer f(1)=31

OpenStudy (freckles):

@kai47 on your previous question you had the following: \[f(n+1)=f(n)-8 \\ \text{ this says term is equal to the previous term minus 8}\] Here you have term=previous term minus 3 you are going backwards in your decreasing sequence so which means you will find the number increasing that is f(3) will be bigger than f(4) f(2) will be bigger than f(3) f(1) will be bigger than f(2) The previous question we just kept subtracting 8 (the common difference was -8) Here your common difference is -3...but you will not be subtract 3 because you are going backwards.. you will adding 3

OpenStudy (freckles):

let's check...

OpenStudy (freckles):

f(4)=22 f(3)=22+3=25 f(2)=25+3=28 f(1)=28+3=31 you did it! :)

OpenStudy (anonymous):

yay I figured it out from your help on my previous question

OpenStudy (freckles):

cool stuff